# On the nature of buckling instability in galactic bars

### v1

@@ 3 First phase: growing the distortion
Figure 3 shows the distortion of the stellar component of the galaxy along the vertical direction ($z$) as it develops in the first stage of buckling. The maps in the subsequent panels show the mean position of the stars along $z$ in different simulation outputs separated by $0.05$ Gyr, starting with $t=4.1$ Gyr. In the following panels the distortion becomes more and more pronounced, preserving its shape with respect to the bar, which is oriented along the $x$ axis in all panels, until about $t=4.5$ Gyr, after which time it starts to wind up. This first stage of buckling can be described in a simple way in terms of a driven harmonic oscillator.
Let us suppose some stellar orbits of the bar are caught at vertical resonance and a small distortion out of the disk plane forms in the bar, like the one shown in the upper left panel of Fig. 3. The other stars will feel a vertical force with a distribution tracing that of the existing distortion. The vertical oscillations of the stars will then be governed by the equation of the driven harmonic oscillator
$\ddot{z}+\omega^{2}z=F_{0}\cos(\omega_{\rm f}t)$ (2)
@@
where $\omega$ is the oscillation frequency of the star in the vertical direction, $\omega_{\rm f}$ is the frequency of the force and $F_{0}$ is its initial amplitude. Let us express the vertical frequencies in terms of the horizontal frequencies, $\omega=f_{z}/f_{x}$. As argued in the previous Section, we can assume that the horizontal frequencies $f_{x}$ do not change in this stage of buckling. From the previous studies we know (L19) that before buckling $42<f_{z}/f_{x}<$ 4; the distribution of these frequency ratios is shown in Fig. 2 in the form of the blue histogram. Expressed in the same way, the force frequency $\omega_{\rm f}=2$.
@@
Figure 8: Pattern speed of the vertical distortion during and after buckling as a function of the amplitude of oscillations along the bar. The left panel shows the approximately constant pattern speed during buckling, assuming the evolved $f_{z}^{\prime}$ and unevolved $f_{x}$. The middle panel shows the strongly decreasing pattern speed after buckling with evolved $f_{z}^{\prime}$ and $f_{x}^{\prime}$. The right panel shows $(f_{x}^{\prime}+f_{\rm p})/3$ which turns out to be equal to the pattern speed after buckling.
The solution of equation (2) with the natural choice of initial conditions $z(0)=z_{0}$ and $\dot{z}(0)=0$ is
$z(t)=z_{0}\cos(\omega t)+\frac{F_{0}}{\omega^{2}-\omega_{\rm f}^{2}}[\cos(\omega_{\rm f}t)-\cos(\omega t)].$ (3)
The well-known outcome of the driven harmonic oscillator is to change the initial frequency $\omega$ toward the frequency of the force, $\omega_{\rm f}$. Figure 5 shows a few examples of orbits that illustrate this transformation. The upper panels of the Figure present the solutions for $z(t)$ for three combinations of $\omega$ and $F_{0}$, $(\omega,F_{0})=$ (2.5, 3), (3, 6) and (3.5, 10) that cover the range of interesting values of $\omega$ found to be present in the bar before buckling (Fig. 2). The values of $F_{0}$ were adjusted to produce sufficient effect on the orbits. In each panel the blue line corresponds to the initial (unforced) oscillation of the star $z(t)=z_{0}\cos(\omega t)$ and the red one the solution under the effect of the force. One can see that the effect of the force is to increase the amplitude of the oscillation, but also to change its frequency. While within $4\pi$ the star oscillates 5, 6 and 7 times for $\omega=2.5,3$ and 3.5, respectively, before the forcing, it makes only 4 oscillations after forcing. The lower panels of Fig. 5 show the corresponding 3D orbits of the stars in the form of the Lissajous curves combining the motion along $z$ with the ellipse or x1-type orbit in the $xy$ plane. The blue and red orbits of the different columns show the 3D motion of the stars before and after the action of the force and the three panels in the lowest row show them together for comparison. While the orbits are not exactly the banana-type orbits because of the variation of the amplitude, they are remarkably similar.
@@
One can now imagine how the stellar orbits in the bar evolve in this phase of buckling. After a small perturbation out of the disk plane appears, the stars with $\omega$ only slightly larger than 2 will have their amplitudes increased, increasing the force. This in turn will affect orbits with higher values of $\omega$ bringing them all to $\omega\sim 2$. This is what we see in the first nine panels of Fig. 3: the distortion grows as more orbits join the resonance preserving its orientation with respect to the bar. This process takes about 0.4 Gyr (from $t=4.1$ till $t=4.5$ Gyr) in this simulation. At $t=4.5$ Gyr many of the stars that were on x1 orbits initially are following the banana-like resonant orbits.
@@
This simple description allows to understand how the distortion can build up producing orbits similar in shape to the well-known banana orbits. Such orbits would produce a distortion that is symmetrical with respect to the $x$ axis of the bar. As one can see from Fig. 3 this is not exactly the case. While the yellow and red part of the distortion, signifying departure above the disk plane, is approximately aligned with the bar ends, the blue part (below the plane) is not exactly perpendicular to the bar. This means that the average orbit of the stars in this stage looks more like the one shown in Fig. 6, which has a minimum shifted with respect to the $y$ axis.
@@
Interestingly, this kind of orbit also explains the mean velocity of stars during this phase of buckling discussed by Li et al. (2023). Figure 7 shows the mean velocity of the stars in the horizontal and vertical direction, $v_{x}$ and $v_{z}$ (upper and lower panel, respectively) at the time of maximum distortion ($t=4.5$ Gyr). The motion of the stars can be understood if the stars on average follow the orbit shown in black. The components of the stellar motion (black arrows along the orbit) along the $x$ and $z$ directions agree with the overall average motion of all stars in these directions shown by the white arrows. This average orbit probably originates from the combination of banana and anti-banana ($\infty$-like) orbits.
@@
The maximum distortion of the bar in the first phase of buckling, at time $t_{\rm b}=4.5$ Gyr can be approximated as $z_{\rm d}(R,\phi,t_{\rm b})=z_{0}(R)\cos m\phi$ with $m=2$. As discussed by Binney & Tremaine (2008) in the context of galactic warps (their section 6.6.1), the evolution of such a distortion can be described as a propagation of a kinematic bending wave with a pattern speed $\omega_{\rm p}=\Omega-\nu/2$ (a different symbol $\omega_{\rm p}$ is used here to distinguish it from the pattern speed of the bar, $\Omega_{\rm p}$ or $f_{\rm p}$ in the notation of this paper). At resonance during buckling and using the notation adopted here, $\omega_{\rm p}=f_{x}+f_{\rm p}-f_{z}^{\prime}/2=f_{\rm p}$ since $f_{z}^{\prime}=2f_{x}$ at this time. This means that the bending wave propagates with the pattern speed of the bar and therefore the distortion pattern seen in Fig. 3 remains stationary in the reference frame of the bar until $t_{\rm b}=4.5$ Gyr. This is confirmed by the distribution of $f_{x}+f_{\rm p}-f_{z}^{\prime}/2$ plotted as a function of the amplitude of the oscillations along the horizontal axis, $a_{x}$, in the left panel of Fig. 8. This combination of frequencies remains approximately constant with radius during buckling. Here, as before, I assumed that the frequencies along the vertical direction have already changed ($f_{z}^{\prime}$), while those along the bar have not yet evolved ($f_{x}$).
@@
Figure 9: Passage of the bending wave. The images from top to bottom illustrate the evolution of the distortion in time at one selected distance from the center of the bar. As the minimum of the curve moves along the ellipse in the $xy$ plane, the distortion changes from the smile-like to the frown-like in the edge-on view. The lowest image shows the combination of all the shapes.
@@
Figure 10: Configuration leading to the distortion seen in the map of average positions of the stars along the vertical direction (middle panel) and in the surface density of the stellar component seen edge-on (along $y$, lower panel) for one selected time, $t=4.85$ Gyr. The dashed arrows indicate the correspondence between the distortions rotated by the bending wave differently at different radii and the maxima and minima of the distributions of the stars in the face-on position map and edge-on density map.
@@
Figure 11: Dependence of frequency of oscillations along the bar on their amplitude before (left panel) and after (right panel) buckling.
@@
Figure 12: Relations between amplitudes of oscillations along the bar (left panel) and their frequencies (right panel) before and after buckling. White dashed lines indicate equality between the measured values and the black points mark the values corresponding to the first four orbits shown in Fig. 13 with properties listed in the first four rows of Table 1.
4 Second phase: winding up the distortion
The evolution of the distortion in the vertical direction changes after $t_{\rm b}=4.5$ Gyr. As demonstrated by the three panels of the lowest row in Fig. 3 ($t=4.55-4.65$ Gyr), the pattern starts to wind up and continues to do so in the later stages illustrated in Fig. 4. The passage of the bending wave at one particular distance from the center of the bar is visualized in Fig. 9 where the distortion of the bar is seen to evolve from the smile-like to the frown-like in the edge-on view. The same happens at all radii, but at a different rate, which leads to a more complicated distortion patterns seen later on in Fig. 4. The connection between the passage of the bending wave and the distortion patterns is illustrated in Fig. 10 for $t=4.85$ Gyr, one of the outputs shown in Fig. 4. The maximum distortion was smile-like at all radii at $t_{\rm b}=4.5$ Gyr. Between $t=4.5$ and 4.85 Gyr the bending wave rotated the distortion differently at different radii, and more so for the radii closer to the center (upper panel). This evolution creates a pattern of maxima and minima in the map of the average positions of the stars (middle panel) along the bar. The same pattern can also be discerned in the edge-on view of the surface density of the stellar component of the galaxy at this time (lower panel). Figure 4 shows that as the winding-up progresses, the pattern of maximal and minimal distortion moves outward toward the end of the bar but is still well visible for example at $t=5.05$ Gyr. Later on the distortion becomes so tightly wound that it ceases to be visible in the inner 5 kpc. The distortion, however, is still present at radii $R\sim 7$ kpc so in a sense buckling continues at these outer radii, as I discuss below.
@@ 4 Second phase: winding up the distortion
Thus, in the second phase of buckling the pattern speed of the bending wave is no longer constant with radius. Indeed, the middle panel of Fig. 8 shows that after buckling $f_{x}^{\prime}+f_{\rm p}-f_{z}^{\prime}/2$ decreases strongly as a function of $a_{x}^{\prime}$ (now all the quantities are primed and measured in the period of time after buckling, as discussed in Section 2). Interestingly, this combination of frequencies turns out to behave almost exactly the same way as $(f_{x}^{\prime}+f_{\rm p})/3$ (or $\Omega^{\prime}/3$ in the standard notation), which is shown in the right panel of Fig. 8. The only difference is that in the middle panel a small remnant of the constant pattern speed remains for 4.5 kpc $<a_{x}^{\prime}<$ 7 kpc. This corresponds to the small peak of the remaining resonant (banana) orbits with $f_{z}^{\prime}/f_{x}^{\prime}=2$ seen in the red histogram of Fig. 2. The equality between these two expressions, $f_{x}^{\prime}+f_{\rm p}-f_{z}^{\prime}/2=(f_{x}^{\prime}+f_{\rm p})/3$ leads immediately to equation (1) after rearranging the terms.
@@
The change in the behavior of the pattern speed of the bending wave after buckling can be traced to the change in the frequency along the bar, since this is the only quantity that varies from $f_{x}$ to $f_{x}^{\prime}$ between the left and middle panels of Fig. 8. Indeed, comparing $f_{x}(a_{x})$ before buckling with $f_{x}^{\prime}(a_{x}^{\prime})$ after buckling in Fig. 11, one can immediately see that the horizontal frequencies increase. This can be further appreciated in Fig. 12 which shows separately how the amplitude of the oscillations and the frequency evolved. This figure demonstrates that after buckling the amplitudes were on average decreased ($a_{x}^{\prime}<a_{x}$) while the frequencies were increased ($f_{x}^{\prime}>f_{x}$). The difference between the frequencies before and after buckling vanishes near the end of the bar, around $R\sim 7$ kpc. In this region $f_{x}^{\prime}+f_{\rm p}-f_{z}^{\prime}/2=(f_{x}^{\prime}+f_{\rm p})/3$ is still equal to the pattern speed of the bar $f_{\rm p}$ because $f_{x}^{\prime}=2f_{\rm p}$, and this is why the distortion survives.

### v2

@@ 3 First phase: Growing the distortion
Figure 3 shows the distortion of the stellar component of the galaxy along the vertical direction ($z$) as it develops in the first stage of buckling. The maps in the subsequent panels show the mean position of the stars along $z$ in different simulation outputs separated by $0.05$ Gyr, starting with $t=4.1$ Gyr. In the following panels, the distortion becomes more and more pronounced, preserving its shape with respect to the bar, which is oriented along the $x$-axis in all panels, until about $t=4.5$ Gyr, after which time it starts to wind up.
The initial stage of the formation of the distortion can be described using the Mathieu equation (Binney, 1981). This formalism predicts instability of stellar orbits with resonant frequencies of $f_{z}=2f_{x}$, which are thus able to increase their oscillations in the $z$ direction. However, as can be seen from Fig. 1 (left panel) and Fig. 14 of L19 (middle left panel), the number of stars with such frequencies is very low before buckling. Although weakly populated, the resonance is by no means narrow, since the stars with resonant frequencies lie in a wide range of radii: between 0.5 and 4 kpc. The mechanism that is able to increase the number of such stars can be described in a simple way in terms of a driven harmonic oscillator.
I now suppose that some stellar orbits of the bar are caught at vertical resonance and that a small distortion out of the disk plane forms in the bar, such as the one shown in the upper left panel of Fig. 3. The other stars will feel a vertical force with a distribution tracing that of the existing distortion. The vertical resonance thus provides a seeding perturbation, which modifies the first orbits, and they change the mass distribution in the bar and its potential. The modified potential then changes the shape of subsequent orbits, which increase the overall distortion creating a kind of feedback loop. The picture involving the forced harmonic oscillator also helps us understand why one type of distortion (smile or frown) is selected and maintained. Once a small distortion appears due to the resonance, it builds up and enhances the same shape, because the force from the distorted mass distribution can only modify the orbits into this particular shape.
The vertical oscillations of the stars can therefore be described by the modified equation of the driven harmonic oscillator of the form
$\ddot{z}+\omega^{2}z=F_{0}(1-\rm{e}^{-\alpha t})\cos(\omega_{\rm f}t),$ (2)
where $\omega$ is the oscillation frequency of the star in the vertical direction, $\omega_{\rm f}$ is the frequency of the force, while $F_{0}$ and $\alpha$ are constants. The classical term $\cos(\omega_{\rm f}t)$ describes the variability of the force due to the distortion of the mass distribution. The additional term $(1-\rm{e}^{-\alpha t})$ represents its growth in time up to the maximum value, $F_{0}$, with the growth rate controlled by $\alpha$.
@@
Let us express the vertical frequencies in terms of the horizontal frequencies, $\omega=f_{z}/f_{x}$. As argued in the previous section, we can assume that the horizontal frequencies, $f_{x},$ do not change in this stage of buckling. From the previous studies (L19), we know that before buckling, $42<f_{z}/f_{x}<$ 4; the distribution of these frequency ratios is shown in Fig. 2 in the form of the blue histogram. Expressed in the same way, the force frequency $\omega_{\rm f}=2$.
@@
Figure 7: Pattern speed of vertical distortion during and after buckling as function of amplitude of oscillations along the bar. The left panel shows the approximately constant pattern speed during buckling, assuming the evolved $f_{z}^{\prime}$ and unevolved $f_{x}$. The middle panel shows the strongly decreasing pattern speed after buckling with evolved $f_{z}^{\prime}$ and $f_{x}^{\prime}$. The right panel shows $(f_{x}^{\prime}+f_{\rm p})/3,$ which turns out to be equal to the pattern speed after buckling.
The solution of Equation (2) with the natural choice of initial conditions $z(0)=z_{0}$ and $\dot{z}(0)=0$ can be found analytically with Mathematica software, but it is too complicated to cite it here in full. Instead, I discuss a few examples. The well-known outcome of the driven harmonic oscillator is to change the initial frequency, $\omega,$ to the frequency of the force, $\omega_{\rm f}$, and it is also the case here. Figure 4 shows three representative orbits that illustrate this transformation. The upper panels of the figure present the solutions for $z(t)$ for $\omega=2.5$, 3, and 3.5, which cover the range of interesting values of $\omega$ found to be present in the bar before buckling (Fig. 2). For all orbits, the values of the constants were assumed to be $F_{0}=10$ and $\alpha=0.1$, which turn out to produce a sufficient effect.
In each panel, the blue line corresponds to the initial (unforced) oscillation of the star $z(t)=z_{0}\cos(\omega t),$ and the red one to the solution under the effect of the force. One can see that the effect of the force is to increase the amplitude of the oscillation (from the initial small value of $z_{0}=0.5$ kpc), but also to change its frequency. While within $4\pi$ the star oscillates five, six, and seven times for $\omega=2.5,3,$ and 3.5, respectively, before the forcing, it only oscillates four times after forcing. It should be noted that the orbits with initial frequencies differing more from the frequency of the force need more time to transform.
The lower panels of Fig. 4 show the corresponding 3D orbits of the stars in the form of the Lissajous curves combining the motion along $z$ with the ellipse in the $xy$ plane. The ellipses are the x1-type orbits, which are well known from classical studies of the orbital structure of bars (Contopoulos & Papayannopoulos, 1980; Skokos et al., 2002). The blue and red orbits of the different columns show the 3D motion of the stars without and with the action of the force, and the three panels in the lowest row show them together for comparison. While the modified orbits are not exactly the banana-type orbits because of the variation of the amplitude, they are remarkably similar.
@@
One can now imagine how the stellar orbits in the bar evolve in this phase of buckling. After a small perturbation out of the disk plane appears, the stars with $\omega$ only slightly larger than two will have their amplitudes strongly increased (left column of Fig. 4), thus increasing the force. This in turn will affect orbits with higher values of $\omega,$ bringing them all to $\omega\sim 2$ (middle and right columns of Fig. 4). This is what we see in the first nine panels of Fig. 3: the distortion grows as more orbits join the resonance preserving its orientation with respect to the bar. This process takes about 0.4 Gyr (from $t=4.1$ till $t=4.5$ Gyr) in this simulation. At $t=4.5$ Gyr, many of the stars that were on x1 orbits initially now follow the banana-like resonant orbits.
@@
This simple description makes it possible to understand how the distortion can build up, creating orbits that are similar in shape to the well-known banana orbits. Such orbits would produce a distortion that is symmetrical with respect to the $x$-axis of the bar. As one can see from Fig. 3, this is not exactly the case. While the yellow and red part of the distortion, signifying departure above the disk plane, is approximately aligned with the bar ends, the blue part (below the plane) is not exactly perpendicular to the bar. This means that the average orbit of the stars at this stage looks more like the one shown in Fig. 5, which has its minima shifted with respect to the $y$-axis.
@@
Interestingly, this kind of orbit also explains the mean velocity of stars during this phase of buckling discussed by Li et al. (2023). Figure 6 shows the mean velocity of the stars in the horizontal and vertical directions, $v_{x}$ and $v_{z}$ (upper and lower panels, respectively) at the time of maximum distortion ($t=4.5$ Gyr). The motion of the stars can be understood if on average the stars follow the orbit shown in black. The components of the stellar motion (black arrows along the orbit) along the $x$ and $z$ directions agree with the overall average motion of all stars in these directions shown by the white arrows. This average orbit probably originates from the combination of banana and anti-banana ($\infty$-like) orbits.
@@
The maximum distortion of the bar in the first phase of buckling, at time $t_{\rm b}=4.5$ Gyr, can be approximated as $z_{\rm d}(R,\phi,t_{\rm b})=z_{0}(R)\cos m\phi$ with $m=2$. As discussed by Binney & Tremaine (2008) in the context of galactic warps (their Section 6.6.1), the evolution of such a distortion can be described as a propagation of a kinematic bending wave with a pattern speed of $\omega_{\rm p}=\Omega-\nu/2$ (a different symbol, $\omega_{\rm p},$ is used here to distinguish it from the pattern speed of the bar, $\Omega_{\rm p},$ or $f_{\rm p}$ in the notation of this paper). At resonance during buckling and using the notation adopted here, $\omega_{\rm p}=f_{x}+f_{\rm p}-f_{z}^{\prime}/2=f_{\rm p,}$ since $f_{z}^{\prime}=2f_{x}$ at this time. This means that the bending wave propagates with the pattern speed of the bar, and therefore the distortion pattern seen in Fig. 3 remains stationary in the reference frame of the bar until $t_{\rm b}=4.5$ Gyr. This is confirmed by the distribution of $f_{x}+f_{\rm p}-f_{z}^{\prime}/2,$ which is plotted as a function of the amplitude of the oscillations along the horizontal axis, $a_{x}$, in the left panel of Fig. 7. This combination of frequencies remains approximately constant with radius during buckling. Here, as before, it is assumed that the frequencies along the vertical direction have already changed ($f_{z}^{\prime}$), while those along the bar have not yet evolved ($f_{x}$).
@@
Figure 8: Passage of bending wave. The images from top to bottom illustrate the evolution of the distortion in time at one selected distance from the center of the bar. As the minimum of the curve moves along the ellipse in the $xy$ plane, the distortion changes from smile-like to frown-like in the edge-on view. The lowest image shows the combination of all the shapes.
@@
Figure 9: Configuration leading to distortion seen in map of average positions of stars along vertical direction (middle panel) and in surface density of stellar component seen edge-on (along $y$, lower panel) for one selected time, $t=4.85$ Gyr. The dashed arrows indicate the correspondence between the distortions rotated by the bending wave differently at different radii and the maxima and minima of the distributions of the stars in the face-on position map and edge-on density map.
@@
Figure 10: Dependence of frequency of oscillations along bar on their amplitude before (left panel) and after (right panel) buckling.
@@
Figure 11: Relations between amplitudes of oscillations along bar (left panel) and their frequencies (right panel) before and after buckling. White dashed lines indicate equality between the measured values and the black points mark the values corresponding to the first four orbits shown in Fig. 12, with properties listed in the first four rows of Table 1.
@@
4 Second phase: Winding up the distortion
The evolution of the distortion in the vertical direction changes after $t_{\rm b}=4.5$ Gyr. As demonstrated by the three panels of the lowest row in Fig. 3 ($t=4.55-4.65$ Gyr), the pattern starts to wind up and continues to do so in the later stages illustrated in Fig. 13 of the appendix. The passage of the bending wave at one particular distance from the center of the bar is visualized in Fig. 8, where the distortion of the bar is seen to evolve from smile-like to frown-like in the edge-on view. The same happens at all radii, but at a different rate, which leads to the more complicated distortion patterns seen at later times in Fig. 13.
The connection between the passage of the bending wave and the distortion patterns is illustrated in Fig. 9 for $t=4.85$ Gyr, one of the outputs shown in Fig. 13. The maximum distortion was smile-like at all radii at $t_{\rm b}=4.5$ Gyr. Between $t=4.5$ and 4.85 Gyr, the bending wave rotated the distortion differently at different radii, and more so for those closer to the center (upper panel). This evolution creates a pattern of maxima and minima in the map of the average positions of the stars (middle panel) along the bar. The same pattern can also be discerned in the edge-on view of the surface density of the stellar component of the galaxy at this time (lower panel).
Figure 13 shows that as the winding-up progresses, the pattern of maximal and minimal distortion moves outward toward the end of the bar but is still clearly visible, for example at $t=5.05$ Gyr. Later on the distortion becomes so tightly wound that it ceases to be visible in the inner 5 kpc. The distortion, however, is still present at radii of $R\sim 7$ kpc, so in a sense buckling continues at these outer radii, as discussed below.
Thus, in the second phase of buckling the pattern speed of the bending wave is no longer constant with radius. Indeed, the middle panel of Fig. 7 shows that after buckling $f_{x}^{\prime}+f_{\rm p}-f_{z}^{\prime}/2$ decreases strongly as a function of $a_{x}^{\prime}$ (now all the quantities are primed and measured in the period of time after buckling, as discussed in Section 2). Interestingly, this combination of frequencies turns out to behave almost exactly the same way as $(f_{x}^{\prime}+f_{\rm p})/3$ (or $\Omega^{\prime}/3$ in the standard notation), which is shown in the right panel of Fig. 7. The only difference is that in the middle panel a small remnant of the constant pattern speed remains for 4.5 kpc $<a_{x}^{\prime}<$ 7 kpc. This corresponds to the small peak of the remaining resonant (banana) orbits with $f_{z}^{\prime}/f_{x}^{\prime}=2$ seen in the red histogram of Fig. 2. The equality between these two expressions, $f_{x}^{\prime}+f_{\rm p}-f_{z}^{\prime}/2=(f_{x}^{\prime}+f_{\rm p})/3,$ immediately leads to Equation (1) after rearrangement of the terms.
@@ 4 Second phase: Winding up the distortion
The change in the behavior of the pattern speed of the bending wave after buckling can be traced to the change in the frequency along the bar, since this is the only quantity that varies from $f_{x}$ to $f_{x}^{\prime}$ between the left and middle panels of Fig. 7. Indeed, by comparing $f_{x}(a_{x})$ before buckling with $f_{x}^{\prime}(a_{x}^{\prime})$ after buckling in Fig. 10, one can immediately see that the horizontal frequencies increase. This can be further appreciated in Fig. 11, which shows separately how the amplitude of the oscillations and the frequency evolved. This figure demonstrates that after buckling the amplitudes were on average decreased ($a_{x}^{\prime}<a_{x}$), while the frequencies were increased ($f_{x}^{\prime}>f_{x}$). The difference between the frequencies before and after buckling vanishes near the end of the bar at around $R\sim 7$ kpc. In this region, $f_{x}^{\prime}+f_{\rm p}-f_{z}^{\prime}/2=(f_{x}^{\prime}+f_{\rm p})/3$ is still equal to the pattern speed of the bar $f_{\rm p}$ because $f_{x}^{\prime}=2f_{\rm p}$, and this is why the distortion survives.